# Pretrain Finite Element Method: A Pretraining and Warm-start Framework for PDEs via Physics-Informed Neural Operators

### v2

@@ Abstract
Transolver , Neural operator , Physics-informed neural operator , Physics-informed neural network , AI for PDEs , Solid Mechanics
@@ 4 Results
Problem Layers Tokens Learning rate Relative error (mean/std) Time: s (PFEM/FEM) Iteration (Speedup: FEM/PFEM) Optimizer Elastic plate (G) 3 64 0.002 0.0113 / 0.0344 0.0257 / 5.117 6.24 = 519.88 / 83.32 Adam Elastic plate (G+M) 4 128 0.002 0.0262 / 0.0183 0.0633 / 5.826 6.43 = 834.84 / 129.69 Adam Elastic plate (G+M+B) 4 128 0.002 0.0209 / 0.0157 0.0694 / 5.984 9.18 = 910.08 / 99.06 Adam Hyperelastic beam (M+B) 4 128 0.002 0.0661 / 0.0453 0.0914 / 36.147 3.66 = 4.90 / 1.34 Adam Hyperelastic Cook (G+M+B) 4 128 0.002 0.0304 / 0.0460 0.122 / 93.752 5.41 = 8.38 / 1.55 Adam
@@ 5.1 Out-of-Distribution Generalization of the PFEM Pretraining Stage
Fig.˜19 illustrates the out-of-distribution generalization performance of the PFEM pretraining stage in the linear elastic plate problem. It can be observed that PFEM is able to accurately predict the displacement field, showing good agreement with the FEM reference solution. Compared with the in-distribution cases, a moderate increase in the prediction error is observed, which is expected under distribution shifts. Nevertheless, the overall solution quality remains high, demonstrating the robustness and strong extrapolation capability of the PFEM pretraining stage.
@@ 5.6 Comparison between PFEM warm-start and preconditioned iterative methods
The common objective of both approaches is to reduce the number of iterations required for convergence. However, traditional preconditioning techniques aim to approximate the inverse of the stiffness matrix in order to improve the conditioning of the linear system, thereby increasing the convergence rate of the iterative solver. In contrast, PFEM accelerates convergence by providing a more accurate initial guess, which effectively changes the starting point of the iteration rather than the spectral properties of the system matrix.
@@ Appendix A Comparison of computational graphs for automatic and explicit differentiation
The upper part of Fig.˜20 illustrates the computational graph associated with automatic differentiation of a neural operator. When constructing differential operators via AD, new computational graphs are generated recursively through the chain rule. Let $C_{T}$ denote the number of basic computational steps required to evaluate the mapping $\boldsymbol{x}\mapsto T(\boldsymbol{x})$ through the neural operator. Then:
@@ Appendix B Iterative algorithms of the finite element method
There are two major numerical approaches to solve Equation˜16: direct methods and iterative methods. For large-scale problems, iterative methods are typically preferred due to their favorable computational complexity. The iterative scheme associated with Equation˜16 can be written as
@@
Substituting the above expressions into Equation˜20 yields
@@ Appendix C Benefit of the warm-start stage in PFEM
Substituting Equations˜41 and 42 into Equation˜43 yields
@@
The benefit brought by a good initial guess becomes less pronounced when the convergence tolerance (tol) is set to be very small. Rearranging Equation˜47 gives
@@
Let $\varepsilon=\|\boldsymbol{B}^{k}\|$, then Equation˜48 can be rewritten as
@@
The rate $R(\boldsymbol{B})$ depends only on the iteration matrix $\boldsymbol{B}$ and is independent of the iteration index $k$ and the initial guess. For sufficiently large $k$, substituting Equation˜51 into Equation˜50 yields
@@
Different choices of the convergence tolerance essentially correspond to different target magnitudes of $\|\boldsymbol{e}^{(k)}\|$. Let $\texttt{tol}=10^{-m}$. In practice, one may regard $\|\boldsymbol{e}^{(k)}\|$ as being on the order of tol at convergence. Then Equation˜49 becomes
@@
Substituting Equation˜53 into Equation˜52 yields
@@
From Equation˜54, the term $m\ln 10/R(\boldsymbol{B})$ characterizes the influence of the convergence tolerance on the iteration number, whereas $\ln\|\boldsymbol{e}^{(0)}\|/R(\boldsymbol{B})$ quantifies the influence of the initial guess. Therefore, as the tolerance becomes tighter (i.e., $m$ increases), the relative contribution of the initial guess to the overall iteration count becomes smaller, implying a diminishing speedup benefit from improved initialization when extremely small tolerances are required.
Appendix D Supplementary code

### v3

@@ Abstract
Transolver , Neural operator , Physics-informed neural operator , Physics-informed neural network , AI for PDEs
@@ 5 Results
Problem Layers Tokens Learning rate Relative error (mean/std) Time: s (PFEM/FEM) Iteration (Speedup: FEM/PFEM) Optimizer Elastic plate (G) 3 64 0.002 0.0113 / 0.0344 0.0257 / 5.117 6.24 = 519.88 / 83.32 Adam Elastic plate (G+M) 4 128 0.002 0.0262 / 0.0183 0.0633 / 5.826 6.43 = 834.84 / 129.69 Adam Elastic plate (G+M+B) 4 128 0.002 0.0209 / 0.0157 0.0694 / 5.984 9.18 = 910.08 / 99.06 Adam Hyperelastic beam (M+B) 4 128 0.002 0.0661 / 0.0453 0.0914 / 36.147 3.66 = 4.90 / 1.34 Adam Hyperelastic Cook (G+M+B) 4 128 0.002 0.0304 / 0.0460 0.122 / 93.752 5.41 = 8.38 / 1.55 Adam 3D Homogenization (Geometry+material) 4 64 0.0005 0.0949/0.0384 0.074/150.0 2.727=40.9/15.0 Adam
@@ 5.2 Nonlinear hyperelasticity
5.3 Three-dimensional solid mechanics homogenization
Homogenization plays a fundamental role in multiscale modeling in solid mechanics [hashin1983analysis, guedes1990preprocessing, hassani1998review, nguyen2025universal], and is closely related to the concept of the representative volume element (RVE). Homogenization aims to determine the effective macroscopic material properties by solving boundary value problems, typically subject to periodic boundary conditions, defined on a RVE. Therefore, homogenization serves as an essential tool for studying multiscale physical phenomena.
Traditional numerical homogenization methods rely heavily on finite element analysis and often require substantial computational resources, particularly for problems involving complex geometries, heterogeneous materials, and high-resolution discretizations. Recently, operator learning approaches have attracted increasing attention for modeling multiscale mechanical behavior in solid mechanics [wang2026pretraining, nguyen2025universal, harandi2025spifol]. The pretraining stage of PFEM employs the Transolver model, which belongs to the class of operator learning. Therefore, in this subsection we investigate the performance of PFEM for three-dimensional homogenization problems in solid mechanics.
Various homogenization frameworks exist in the literature. Here, we adopt the classical numerical homogenization approach proposed by Andreassen et al. [andreassen2014determine], which aims to compute the effective macroscopic elastic tensor of periodic composite materials. The detailed homogenization formulation is in [wang2026pretraining]. In [wang2026pretraining], the Fourier neural operator (FNO) was trained in a data-driven manner. However, our manuscript trains the Transolver purely using physics-based constraints without data, following the PFEM framework proposed in this study.
The loss function is derived from the minimum potential energy principle of linear elasticity:
$\mathcal{L}$ $=\frac{1}{6}\sum_{IJ}^{6}\Pi^{(IJ)},$ (15)
$\Pi^{(IJ)}$ $=\int_{\Omega}\frac{1}{2}E_{klpq}\varepsilon_{kl}(\boldsymbol{X}^{(v)(IJ)})\varepsilon_{pq}(\boldsymbol{X}^{(v)(IJ)})\,d\Omega-\int_{\Omega}E_{klpq}\varepsilon_{kl}^{(0)(IJ)}\,d\Omega,$
where $\boldsymbol{X}^{(v)}$ denotes the fluctuation displacement field, which is the target output of the neural operator. $\boldsymbol{\varepsilon}^{(0)(IJ)}$ represents the six prescribed macroscopic loading cases used in the homogenization procedure, corresponding to the six unit strain fields. Further details of the homogenization formulation are provided in D.
The dataset consists of three-dimensional TPMS structures with a resolution of $64\times 64\times 64$. Six types of TPMS geometries are considered: Schoen Gyroid (“solid networks” and “sheet networks”), Schwarz Diamond (“solid networks” and “sheet networks”), and Fischer Koch S (“solid networks” and “sheet networks”), as illustrated in Fig.˜19. For each TPMS type, 600 structures are generated, resulting in a total of 3600 samples. The total dataset size is approximately 103 GB and is publicly available at https://1drv.ms/f/c/f999b24db97d074b/UgBLB325TbKZIID5YoUEAAAAAK57y3r7MVu56Ds.
Both the geometry and the material distributions are randomly generated. The geometric volume fractions range from 0.26 to 0.66, while Poisson’s ratios vary between 0.1 and 0.4. Details of the geometric volume fraction generation are described in E. Both parameters follow uniform distributions. Since the present study focuses on linear elasticity homogenization, Young’s modulus is set to unity during training, which simplifies the computation of the effective elastic tensor through linear scaling. Among the 3600 samples, 3000 are used for training and 600 for testing. It is worth noting that the $64^{3}$ resolution grid represents the entire cubic domain, including both material and void regions. However, when training the Transolver, only the point cloud corresponding to material points is used. As a result, the number of input points varies across different samples.
Figure 19: Six different TPMS structures: (a) Schoen Gyroid: “solid networks”; (b) Schoen Gyroid: “sheet networks”; (c) Schwarz Diamond: “solid networks”; (d) Schwarz Diamond: “sheet networks”; (e) Fischer Koch S: “solid networks”; (f) Fischer Koch S: “sheet networks”.
The architectural details and training configurations of the Transolver are summarized in Table˜1. Fig.˜20 and Fig.˜21 present the prediction results of PFEM, where the finite element solution serves as the reference. The FEM simulations employ $64\times 64\times 64=262{,}144$ linear elements, and the average computational time for a single FEM sample is approximately 150 s.
The results show that PFEM in the pretraining stage predicts displacement fields that agree closely with the FEM reference solutions, with overall relative errors below $10\%$, as summarized in Table˜1. For clarity, only the absolute displacement field corresponding to the first loading case is shown. In practice, six loading cases are required in the homogenization procedure, as detailed in D.
The input to the Transolver consists of four features: the three spatial coordinates and the Poisson’s ratio. The output dimension is 18, corresponding to three displacement components under six loading cases. Notably, PFEM only requires the coordinates of material points in the TPMS structures as input. This differs from FNO-based approaches, which typically require dense 3D grid representations where void regions are filled with zeros [wang2026pretraining]. Consequently, PFEM benefits from the point-cloud-based representation of Transolver, which provides a natural advantage for handling complex geometries. Specifically, the input of PFEM consists of the geometry and material information $\boldsymbol{G},\boldsymbol{M}=\{\boldsymbol{X}^{(i)}\}_{i=1}^{N}$, where $\boldsymbol{X}^{(i)}=\{x_{j}^{(i)},y_{j}^{(i)},z_{j}^{(i)},\nu_{j}^{(i)}\}_{j}^{N^{(i)}}$. The output is the displacement at the corresponding coordinate points $\boldsymbol{U}=\{\boldsymbol{u}^{(i)(1)},\boldsymbol{u}^{(i)(2)},\boldsymbol{u}^{(i)(3)},\boldsymbol{u}^{(i)(4)},\boldsymbol{u}^{(i)(5)},\boldsymbol{u}^{(i)(6)}\}_{i=1}^{N}$, where $\boldsymbol{u}^{(i)(M)}=\{u_{j}^{(i)(M)},v_{j}^{(i)(M)},w_{j}^{(i)(M)}\}_{j}^{N^{(i)}}$, and $M=1,\cdots,6$ corresponds to the six loading cases in the homogenization process.
After obtaining the displacement fields, the effective elastic tensor of size $6\times 6$ is computed, as shown in the last column of Fig.˜20 and Fig.˜21. PFEM in the pretraining stage demonstrates excellent predictive capability for the effective elastic tensor, with maximum relative errors below $3\%$. Meanwhile, PFEM achieves significantly higher computational efficiency than conventional FEM methods, as summarized in Table˜1.
Figure 20: Absolute displacement predictions of PFEM in the pretraining stage for TPMS “solid network” structures: (a) Schoen Gyroid, (b) Schwarz Diamond, (c) Fischer Koch S. The first column shows the FEM reference solution, the second column shows PFEM predictions, the third column presents the error fields, and the last column reports the relative error of the effective elastic tensor.
Figure 21: Absolute displacement predictions of PFEM in the pretraining stage for TPMS “sheet network” structures: (a) Schoen Gyroid, (b) Schwarz Diamond, (c) Fischer Koch S. The columns correspond to FEM reference solutions, PFEM predictions, error fields, and relative errors of the effective elastic tensor, respectively.
The testing errors of PFEM in the pretraining stage are shown in Fig.˜22, where the training process spans 300 epochs. The smallest errors are observed in the displacement component aligned with the principal loading direction. This behavior arises because the energy-based loss function assigns different sensitivities to displacement components under different loading cases. For example, under tensile loading in the $x$ direction, as shown in Fig.˜22a, the displacement error in the $x$ direction is smaller than those in the $y$ and $z$ directions. This is because the $x$-direction displacement is dominant under $x$-direction tensile loading, resulting in larger displacement magnitudes compared with the $y$ and $z$ directions. Consequently, reducing the error in the $x$-direction displacement contributes more significantly to the reduction of the total energy. As a result, during the optimization process PFEM tends to focus more on improving the accuracy of the $x$-direction displacement field.
Figure 22: Testing errors of PFEM in the pretraining stage. (a–f) correspond to tensile loading in the $x$, $y$, and $z$ directions, and shear loading in the xyxy, xzxz, and $yz$ directions. For each loading case, the relative errors of the three displacement components are reported.
Next, we investigate the performance gain of PFEM in the warm-start stage. Fig.˜23 presents the warm-start results for the three-dimensional homogenization problem. In the conventional FEM baseline, the zero vector is used as the initial guess. In contrast, PFEM uses the displacement field predicted by the Transolver as the initial guess for the iterative solver.
It can be clearly observed that PFEM significantly reduces the number of iterations required for convergence, regardless of whether the iterative solver in the warm-start stage is the conjugate gradient (CG) method or the preconditioned conjugate gradient (PCG) method. Compared with standard FEM, PFEM exhibits a more robust and smoother iterative convergence process. As shown in Fig.˜23(a) and (c), the convergence rate is determined solely by the chosen iterative algorithm. Therefore, when the same iterative algorithm is used for both PFEM and FEM, the convergence rates remain identical. The reduction in iteration numbers achieved by PFEM originates entirely from the improved initial guess.
Figure 23: Warm-start results of PFEM for the 3D homogenization problem. The first and second rows correspond to warm-start iterations using the conjugate gradient (CG) and preconditioned conjugate gradient (PCG) methods, respectively. The first column shows the convergence histories of PFEM and FEM, while the second column compares the number of iterations required by PFEM and FEM.
Overall, this subsection demonstrates that PFEM generalizes effectively to complex three-dimensional homogenization problems involving diverse geometries and material distributions. Notably, no special preprocessing is required, and the model only takes point-cloud inputs corresponding to material points. Furthermore, when combined with the warm-start stage, PFEM substantially reduces the number of iterations required to solve large-scale 3D homogenization problems.
@@ 6.1 Out-of-Distribution Generalization of the PFEM Pretraining Stage
Fig.˜24 illustrates the out-of-distribution generalization performance of the PFEM pretraining stage in the linear elastic plate problem. It can be observed that PFEM is able to accurately predict the displacement field, showing good agreement with the FEM reference solution. Compared with the in-distribution cases, a moderate increase in the prediction error is observed, which is expected under distribution shifts. Nevertheless, the overall solution quality remains high, demonstrating the robustness and strong extrapolation capability of the PFEM pretraining stage.
@@ 6.6 Comparison between PFEM warm-start and preconditioned iterative methods
The common objective of both approaches is to reduce the number of iterations required for convergence. However, traditional preconditioning techniques aim to approximate the inverse of the stiffness matrix in order to improve the conditioning of the linear system, thereby increasing the convergence rate of the iterative solver. In contrast, PFEM as illustrated in Fig.˜23 accelerates convergence by providing a more accurate initial guess, which effectively changes the starting point of the iteration rather than the spectral properties of the system matrix.
@@ Appendix A Comparison of computational graphs for automatic and explicit differentiation
The upper part of Fig.˜25 illustrates the computational graph associated with automatic differentiation of a neural operator. When constructing differential operators via AD, new computational graphs are generated recursively through the chain rule. Let $C_{T}$ denote the number of basic computational steps required to evaluate the mapping $\boldsymbol{x}\mapsto T(\boldsymbol{x})$ through the neural operator. Then:
@@ Appendix B Iterative algorithms of the finite element method
There are two major numerical approaches to solve Equation˜17: direct methods and iterative methods. For large-scale problems, iterative methods are typically preferred due to their favorable computational complexity. The iterative scheme associated with Equation˜17 can be written as
@@
Substituting the above expressions into Equation˜21 yields
@@ Appendix C Benefit of the warm-start stage in PFEM
Substituting Equations˜42 and 43 into Equation˜44 yields
@@
The benefit brought by a good initial guess becomes less pronounced when the convergence tolerance (tol) is set to be very small. Rearranging Equation˜48 gives
@@
Let $\varepsilon=\|\boldsymbol{B}^{k}\|$, then Equation˜49 can be rewritten as
@@
The rate $R(\boldsymbol{B})$ depends only on the iteration matrix $\boldsymbol{B}$ and is independent of the iteration index $k$ and the initial guess. For sufficiently large $k$, substituting Equation˜52 into Equation˜51 yields
@@
Different choices of the convergence tolerance essentially correspond to different target magnitudes of $\|\boldsymbol{e}^{(k)}\|$. Let $\texttt{tol}=10^{-m}$. In practice, one may regard $\|\boldsymbol{e}^{(k)}\|$ as being on the order of tol at convergence. Then Equation˜50 becomes
@@
Substituting Equation˜54 into Equation˜53 yields
@@
From Equation˜55, the term $m\ln 10/R(\boldsymbol{B})$ characterizes the influence of the convergence tolerance on the iteration number, whereas $\ln\|\boldsymbol{e}^{(0)}\|/R(\boldsymbol{B})$ quantifies the influence of the initial guess. Therefore, as the tolerance becomes tighter (i.e., $m$ increases), the relative contribution of the initial guess to the overall iteration count becomes smaller, implying a diminishing speedup benefit from improved initialization when extremely small tolerances are required.
Appendix D Numerical homogenization theory
In this section, we briefly introduce the numerical homogenization framework proposed by Andreassen et al. [andreassen2014determine], which aims to compute the effective macroscopic elastic tensor of periodic composite materials. For simplicity, we consider a two-phase material composed of solid and void regions. The elastic tensor of the solid phase is denoted by $E_{ijkl}^{S}$.
To compute the homogenized elastic tensor $E_{IJKL}^{H}$, six independent unit strain fields are applied to the unit cell, including three normal strains and three shear strains. The six macroscopic unit strain tensors $\varepsilon_{pq}^{(0)(IJ)}$ are defined as
$\boldsymbol{\varepsilon}^{(0)(11)}$ $=\begin{bmatrix}1&0&0\\ 0&0&0\\ 0&0&0\end{bmatrix},\quad\boldsymbol{\varepsilon}^{(0)(22)}=\begin{bmatrix}0&0&0\\ 0&1&0\\ 0&0&0\end{bmatrix},\quad\boldsymbol{\varepsilon}^{(0)(33)}=\begin{bmatrix}0&0&0\\ 0&0&0\\ 0&0&1\end{bmatrix}$, (56)
$\boldsymbol{\varepsilon}^{(0)(12)}$ $=\begin{bmatrix}0&1&0\\ 1&0&0\\ 0&0&0\end{bmatrix},\quad\boldsymbol{\varepsilon}^{(0)(13)}=\begin{bmatrix}0&0&1\\ 0&0&0\\ 1&0&0\end{bmatrix},\quad\boldsymbol{\varepsilon}^{(0)(23)}=\begin{bmatrix}0&0&0\\ 0&0&1\\ 0&1&0\end{bmatrix}$.
Here $\varepsilon_{pq}^{(0)(IJ)}$ denotes a macroscopic unit strain field imposed on the unit cell domain $\Omega$. The capital indices $(IJ)$ correspond to the six loading cases and do not follow the Einstein summation convention. The six loading configurations are illustrated in Fig.˜26.
Figure 26: Schematic illustration of numerical homogenization. Six macroscopic unit strain fields are applied to the unit cell, producing the corresponding displacement fields. The homogenized elastic tensor is then computed through the homogenization formula based on the resulting displacement fields.
The homogenized elastic tensor $E_{IJKL}^{H}$ is computed as
$E_{IJKL}^{H}=\frac{1}{|\Omega|}\int_{\Omega}E_{pqrs}\left(\varepsilon_{pq}^{(0)(IJ)}-\varepsilon_{pq}^{(v)(IJ)}\right)\left(\varepsilon_{rs}^{(0)(KL)}-\varepsilon_{rs}^{(v)(KL)}\right)\,d\Omega,$ (57)
where $E_{pqrs}$ denotes the local elastic tensor of the TPMS unit cell,
$E_{ijkl}=\begin{cases}E_{ijkl}^{S},&\text{in the solid phase},\\ 0,&\text{in the void phase}.\end{cases}$ (58)
where $|\Omega|$ denotes the volume of the unit cell. The fluctuation strain field $\varepsilon_{pq}^{(v)(IJ)}$ is defined as
$\varepsilon_{pq}^{(v)(IJ)}=\varepsilon_{pq}\big(\boldsymbol{X}^{(v)(IJ)}\big)=\frac{1}{2}\left(X_{p,q}^{(v)(IJ)}+X_{q,p}^{(v)(IJ)}\right),$ (59)
where $\boldsymbol{X}^{(v)(IJ)}$ denotes the fluctuation displacement field induced by the imposed macroscopic unit strain field. The total displacement field can therefore be written as
$\boldsymbol{X}^{(IJ)}=\boldsymbol{X}^{(0)(IJ)}-\boldsymbol{X}^{(v)(IJ)},$ (60)
where $\boldsymbol{X}^{(0)(IJ)}$ represents the macroscopic affine displacement field corresponding to the uniform strain field, while $\boldsymbol{X}^{(v)(IJ)}$ represents the fluctuation displacement caused by material heterogeneity.
We consider the total potential energy of the unit cell
$\Pi^{(IJ)}$ $=\int_{\Omega}\frac{1}{2}E_{klpq}\varepsilon_{kl}^{(IJ)}\varepsilon_{pq}^{(IJ)}\,d\Omega,$ (61)
$\varepsilon_{kl}^{(IJ)}$ $=\frac{1}{2}\left(X_{k,l}^{(IJ)}+X_{l,k}^{(IJ)}\right)=\varepsilon_{kl}^{(0)(IJ)}-\varepsilon_{kl}^{(v)(IJ)}.$
Taking the first variation of $\Pi^{(IJ)}$ and using the symmetry $E_{klpq}=E_{pqkl}$ yields
$\delta\Pi^{(IJ)}=\int_{\Omega}E_{klpq}\varepsilon_{kl}^{(IJ)}\delta\varepsilon_{pq}^{(IJ)}\,d\Omega=-\int_{\Omega}E_{klpq}\big(\varepsilon_{kl}^{(0)(IJ)}-\varepsilon_{kl}^{(v)(IJ)}\big)\delta\varepsilon_{pq}^{(v)(IJ)}\,d\Omega.$ (62)
By enforcing $\delta\Pi^{(IJ)}=0$, the fluctuation displacement field satisfies the following weak form of the equilibrium equation:
$\int_{\Omega}E_{klpq}\varepsilon_{kl}(\boldsymbol{X}^{(v)(IJ)*})\varepsilon_{pq}(\boldsymbol{X}^{(v)(IJ)})\,d\Omega=\int_{\Omega}E_{klpq}\varepsilon_{kl}(\boldsymbol{X}^{(v)(IJ)*})\varepsilon_{pq}^{(0)(IJ)}\,d\Omega,$ (63)
where $\boldsymbol{X}^{(v)(IJ)*}$ denotes a virtual fluctuation displacement field. In practice, the fluctuation displacement field $\boldsymbol{X}^{(v)(IJ)}$ is typically solved using the finite element method. The fluctuation strain $\boldsymbol{\varepsilon}^{(v)(IJ)}$ can then be computed and substituted into Equation˜57 to obtain the homogenized elastic tensor.
To ensure a unique solution and correctly represent the periodic microstructure, the weak form (63) is solved subject to periodic boundary conditions (PBCs). These conditions enforce identical fluctuation displacements on opposite faces of the unit cell. Assuming $\Omega=[0,L]^{3}$, the PBCs can be expressed as
$\boldsymbol{X}^{(v)(IJ)}(0,x_{2},x_{3})$ $=\boldsymbol{X}^{(v)(IJ)}(L,x_{2},x_{3}),$ (64)
$\boldsymbol{X}^{(v)(IJ)}(x_{1},0,x_{3})$ $=\boldsymbol{X}^{(v)(IJ)}(x_{1},L,x_{3}),$
$\boldsymbol{X}^{(v)(IJ)}(x_{1},x_{2},0)$ $=\boldsymbol{X}^{(v)(IJ)}(x_{1},x_{2},L).$
In conventional finite element implementations, the PBCs are enforced by constraining the degrees of freedom of corresponding nodes on opposite faces of the unit cell.
In PFEM, periodic constraints are imposed through a ring-type coupling strategy. Specifically, pairs of corresponding nodes on opposite faces are connected by virtual elements, and the associated virtual strain energy is incorporated into the overall loss function. It should be emphasized that this virtual strain energy formulation is essentially equivalent to a penalty method, where the stiffness matrix of the virtual elements acts as the penalty parameter controlling the strength of the periodic constraint.
The fluctuation displacement field $\boldsymbol{X}^{(v)(IJ)}$ must be solved for all six loading cases during the homogenization process. In traditional approaches, this requires repeated finite element simulations, which can be computationally expensive, especially when different geometries and material distributions are considered. Therefore, accelerating the computation of displacement fields directly improves the overall efficiency of the homogenization process. This observation motivates the use of PFEM to accelerate numerical homogenization.
Appendix E Generation of TPMS geometries
In this section, we briefly introduce the mathematical description of triply periodic minimal surfaces (TPMS). A minimal surface is characterized by zero mean curvature at every point. TPMS are periodic minimal surfaces that extend infinitely in three-dimensional space. A common way to represent TPMS is through a level-set formulation. In this framework, the geometry is described by a scalar field combined with trigonometric functions, which enables flexible control of both topology and mechanical properties. The level-set equation is generally written as
$\phi(x,y,z)=c,$ (65)
where $x$, $y$, and $z$ denote the spatial coordinates. The scalar field $\phi(x,y,z)$ defines a continuous field in the three-dimensional domain, and the level-set value $c$ partitions a TPMS unit cell into two connected subdomains, corresponding to the solid phase and the void phase, respectively. By varying $c$, the porosity, or equivalently the volume fraction, of the TPMS unit cell can be adjusted. The TPMS surface is thus represented by the iso-surface associated with the level-set value $c$.
One widely used TPMS level-set model is the Schwarz Primitive surface, which is defined as
$\cos(2\pi x)+\cos(2\pi y)+\cos(2\pi z)=c.$ (66)
This level-set equation represents a zero-thickness shell-like minimal surface. To construct load-bearing TPMS lattice materials from such minimal surfaces, two common types of TPMS architectures are considered.
The first type is referred to as solid networks, defined by
$\phi(x,y,z)>c.$ (67)
For solid-network TPMS cells, the region satisfying $\phi(x,y,z)>c$ is identified as the solid phase, while the remaining region is treated as void.
The second type is referred to as sheet networks, defined by
$c\leq\phi(x,y,z)\leq c.$ (68)
As the name suggests, sheet-network TPMS selects a thin volumetric layer around the zero-thickness minimal surface, thereby producing a sheet-like topology within the unit cell. Solid-network and sheet-network TPMS exhibit distinct topological characteristics and effective elastic properties, providing greater flexibility for topology-driven material design.
Another important topological parameter of TPMS is the volume fraction, also known as the relative density. The volume fraction is defined as the ratio of the solid volume to the total volume of the TPMS unit cell. This parameter directly influences the effective elastic tensor and stiffness of the unit cell, and therefore plays a key role in the homogenized elastic response.
In this study, we generated a series of TPMS unit cells to train and validate the proposed neural-operator-based homogenization framework. Three representative TPMS families were considered, namely the Schoen Gyroid, Schwarz Diamond, and Fischer Koch S. Their corresponding level-set equations are given by
$\text{Schoen Gyroid:}\quad\sin(2\pi x)\cos(2\pi y)+\sin(2\pi y)\cos(2\pi z)+\sin(2\pi z)\cos(2\pi x)=c,$ (69)
$\text{Schwarz Diamond:}\quad\cos(2\pi x)\cos(2\pi y)\cos(2\pi z)-\sin(2\pi x)\sin(2\pi y)\sin(2\pi z)=c,$
$\text{Fischer Koch S:}\quad\cos(4\pi x)\sin(2\pi y)\cos(2\pi z)+\cos(2\pi x)\cos(4\pi y)\sin(2\pi z)$
$\qquad\qquad\qquad\quad+\sin(2\pi x)\cos(2\pi y)\cos(4\pi z)=c.$
Appendix F Supplementary code